### v2

# Conan: A Chunkwise Online Network for Zero-Shot Adaptive Voice Conversion

Abstract
Zero-shot online voice conversion (VC) holds significant promise for real-time communications and entertainment. However, current VC models struggle to preserve semantic fidelity under real-time constraints, deliver natural-sounding conversions, and adapt effectively to unseen speaker characteristics. To address these challenges, we introduce Conan, a chunkwise online zero-shot voice conversion model that preserves the content of the source while matching the voice timbre and styles of reference speech. Conan comprises three core components: 1) a Stream Content Extractor that leverages Emformer for low-latency streaming content encoding; 2) an Adaptive Style Encoder that extracts fine-grained stylistic features from reference speech for enhanced style adaptation; 3) a Causal Shuffle Vocoder that implements a fully causal HiFiGAN using a pixel-shuffle mechanism. Experimental evaluations demonstrate that Conan outperforms baseline models in subjective and objective metrics. Audio samples can be found at https://aaronz345.github.io/ConanDemo.
Index Terms:
Voice conversion, Real-time speech processing, Zero-shot learning, Streaming inference
I Introduction
Zero-shot online voice conversion (VC) aims to extract the linguistic content from source speech and, in real time, modify its timbre and styles to match those of a reference speech [1]. This field has shown considerable potential for applications in real-time communication and interactive entertainment.[2] Existing VC approaches typically employ pre-trained models to extract content information from the source speech and combine it with speaker identity embeddings from the reference speech to perform chunkwise online conversion[3].
The field of VC has seen substantial advances [4]. In particular, zero‐shot VC, which aims to transfer out-of-domain speaker identity from reference speech to generated speech, has been explored in recent years [5]. Moreover, real‐time VC systems capable of operating online have been successfully deployed in low‐latency scenarios [6]. However, with the recent growing demand for customizable, high‐quality audio experiences, there is an emerging trend toward zero‐shot online VC that delivers high semantic fidelity, adaptive style transfer, and high-level naturalness [1]. For chunkwise online source speech input, it is essential not only to maintain both content extraction and speech synthesis quality under strict online latency constraints but also to transfer the timbre and styles of unseen reference speakers. Such speech styles encompass emotion, prosody, and other expressive attributes [7]. Adapting these out‐of‐domain timbres and styles to arbitrary source content in VC remains a challenging open problem.
Currently, zero-shot online VC faces three major challenges:
  • Content extraction via chunkwise streaming often compromises quality. Existing VC methods typically employ pretrained feature-extraction models to derive linguistic representations, often leveraging self-supervised models such as HuBERT [8] or WavLM [9]. However, the large receptive fields of these pretrained encoders make them unsuitable for online scenarios with minimal or zero lookahead. Attempts to train a lightweight causal content encoder via distillation may have fallen short in efficiency and accuracy [3]. Alternatively, some approaches rely on phonetic extraction based on automatic speech recognition [10], which incurs additional latency.
  • Achieving adaptive style transfer in zero-shot scenarios is still challenging. Current models often rely on the strong assumption that the reference voice is accessible [11] for model adaptation, which does not always hold in practice. Moreover, many VC approaches focus solely on timbre transfer by directly applying the source speech’s F0 and energy, which contradicts the goal of capturing the reference speaker’s styles [3]. The handful of zero-shot methods that do attempt to generalize to unseen speakers still struggle to balance high speaker similarity with naturalness [12]. This shortfall arises from insufficient modeling of speaker styles, preventing effective alignment of speaker styles with source content [6].
  • Maintaining high quality and naturalness in real-time VC remains difficult. In online VC, the system must process source speech chunk by chunk and synthesize generated speech incrementally, which prevents the direct use of conventional VC models and leads to poor naturalness at chunk boundaries [5]. Leveraging causality with a constrained look-ahead window is a common approach to address this issue. Many previous methods introduce auto-regressive decoders to enforce temporal causality, but this comes at the cost of increased latency [1]. Additionally, streaming models often convert non-causal vocoders into causal ones by zero-padding, introducing spectral artifacts that degrade quality and efficiency [13].
To address these challenges above, we introduce Conan, a chunkwise online network for zero-shot adaptive voice conversion. In zero‐shot online scenarios, Conan preserves the high‐fidelity content of the source speech while matching the timbre and styles of unseen reference speakers, generating highly natural speech on a chunkwise basis. To achieve high-quality chunkwise streaming content extraction, we design the Stream Content Extractor, built on an Emformer architecture and trained using content representations extracted offline by HuBERT. To enable adaptive style transfer, we introduce the Adaptive Style Encoder, which employs clustering-based vector quantization to capture fine-grained style attributes from reference speech and uses an alignment attention mechanism to fuse them with content and timbre information. For high-quality, natural streaming speech synthesis, we adopt a causal-convolution mel decoder and propose the Causal Shuffle Vocoder, a fully causal HiFiGAN that leverages a pixel-shuffle mechanism to eliminate checkerboard artifacts while improving computational efficiency. Experimental evaluations show that Conan outperforms baseline models in real-time content accuracy, conversion quality, and speaker similarity. Conan can achieve a latency as low as 37 ms for the conversion on a single A100 GPU without any engineering optimizations. Overall, our main contributions can be summarized as:
  • We present Conan, a chunkwise online zero-shot voice conversion model that preserves content of source speech while matching timbre and styles from reference speech.
  • We introduce the Stream Content Extractor that leverages Emformer for low-latency streaming content encoding.
  • We design the Adaptive Style Encoder that extracts and aligns styles from reference speech for style adaptation.
  • We present the Causal Shuffle Vocoder that implements a fully causal HiFiGAN using a pixel-shuffle mechanism.
  • Experimental evaluations demonstrate that Conan outperforms baseline models in real-time content accuracy, conversion quality, and speaker similarity.
II Related Works
II-A Zero-Shot Voice Conversion
In practical scenarios, obtaining reference speech during training is often impractical. Therefore, zero-shot VC models are essential to capture and control timbre and styles by disentangling content from speaker characteristics. Chou et al. employ instance normalization to separate speaker and content [14], while Wang et al. retrieve variable‐length speaker embeddings along both temporal and channel dimensions under the guidance of a pre‐trained speaker‐verification model [15]. Ebbers et al. leverage adversarial contrastive predictive coding for fully unsupervised separation of content and speaker [16]. AutoVC squeezes out speaker timbre from content embeddings using an information bottleneck [5], and VQMIVC encodes content via vector quantization, applying mutual information constraints to decorrelate speech components [4]. NANSY trains in a fully self‐supervised manner using wav2vec features alongside Yingram [17], and LM-VC tokenizes speech into semantic tokens via HuBERT and acoustic tokens via SoundStream [18]. Although StreamVoice further advances real‐time separation and recombination of speaker and content using a language model and ASR [1], current online VC systems still exhibit considerable room for improvement in migrating timbre and styles for unseen reference speakers and in accurately extracting linguistic content.
II-B Online Voice Conversion
In online VC, the system must process source speech chunk by chunk and synthesize output incrementally, without access to future context. This often degrades naturalness at chunk boundaries. For streaming applications, causal processing is a critical design consideration. Hayashi et al. [19] propose a streamable version of the non‐autoregressive sequence‐to‐sequence VC model based on FastSpeech2 [20] and NAR‐S2S‐VC [21], incorporating causal convolutions and self‐attention with causal masks. IBF-VC [22] uses the Intermediate bottleneck features (IBFs) to replace Phonetic posteriorgrams (PPGs) in the ASR encoder to capture more fine-grained prosody information, and applies non-streaming teacher guidance for the timbre leakage problem. Yang et al. [6] adapt the originally offline VQMIVC model [4] for real‐time, chunk‐by‐chunk processing. ALO‐VC [23] assembles a streaming system composed of a speaker‐verification model, a streamable phonetic‐posteriorgram extractor, and an F0 extractor. StreamVC [3] demonstrates that a lightweight causal convolutional network can effectively capture soft speech‐unit representations, while StreamVoice [1] introduces autoregressive decoders to enforce temporal causality. Cui et al. [24] convert non‐causal vocoders [25] into causal ones via zero‐padding. However, methods relying on autoregressive structures incur additional latency, and zero‐padded causal vocoders suffer from artifacts and reduced efficiency, limiting overall naturalness. Moreover, none of these approaches achieves robust style adaptation in zero-shot scenarios.
II-C Style Modeling
Modeling speaking styles remains a central challenge in speech research. Prior approaches have largely relied on pre-trained models to capture only a limited set of styles, like wav2vec 2.0 [26], HuBERT [27], and WavLM [9] Attentron [28] introduces an attention mechanism to extract speaker styles from reference samples. ZSM-SS [29] proposes a Transformer-based architecture with an external speaker encoder based on wav2vec 2.0. Daft-Exprt [30] employs a gradient reversal layer to improve reference speaker fidelity in style transfer. GenerSpeech introduces both global and local style adaptors to capture diverse speaking styles [31], while Styler decomposes style into multiple levels of supervision [32]. Yang et al. [6] jointly models speaker identity and global prosody using a GST-based style token network. Mega-TTS 2 employs vector quantization for prosody encoding combined with a language model for prosody transfer [33], and NaturalSpeech 3 uses factorized vector quantization to disentangle prosodic features [34]. CosyVoice integrates x-vectors into a large language model to both disentangle and model prosody [35]. Most of these methods focus on offline speech synthesis, while our work address the challenge of zero-shot online voice conversion by modeling rich speaking styles and maintaining precise content alignment.
Figure 1: The overall architecture of Conan. Online source speech is fed into the system in discrete chunks, and the synthesized output is likewise produced on a chunkwise basis. In Figure (b), the size of the right‐context chunk R is configurable; when it is set to zero, the model operates in a chunkwise causal mode.
III Method
III-A Overview
The architecture of Conan is shown in Figure 1 (a). source speech $y_{sr}$ is input online and fed into the system in chunks, $y_{sr-i}$ represents the i-th chunk of the online source speech. In the Stream Content Extractor, it is extracted as content labels $z_{c-i}\in\mathbb{R}^{T//20}$, where $T$ is the duration of each chunk (ms). Since the offline HuBERT content representation format is used, each content label corresponds to a 20 ms WAV segment. The complete unseen reference speech $y_{rf}$ is passed through a mel extractor to obtain reference mel $m_{rf}$, which is then input to the timbre encoder to produce a one-dimensional global speaker timbre embedding $z_{t}$. Additionally, $m_{rf}$ is also fed into the Adaptive Style Encoder to extract chunk-level style representations encompassing fine-grained emotions, prosody, and other styles. By using aligned attention with $z_{t}$ and $z_{c-i}$, we obtain a style representation $z_{s-i}$ that aligns with the content scale. Next, the causal pitch predictor estimates F0, as explicit pitch modeling greatly aids in conveying speech style [36]. The causal mel decoder then decodes to produce the generated mel $m_{gn-i}$, which is converted into generated speech $y_{gn-i}$ with high quality and stability via a causal shuffle vocoder. This generated speech contains the content of $y_{sr-i}$, the timbre of $y_{rf}$, and the style aligned with the content. The processing time for each chunk is shorter than the chunk’s duration, enabling chunkwise, streaming, low-latency zero-shot online voice conversion. Note that, except for the Stream Content Extractor’s chunk-input-based design, all other components are frame-level causal models, allowing great flexibility in handling diverse chunk size settings.
III-B Stream Content Extractor
Content accuracy is a critical determinant of VC system quality. However, existing approaches based on pre-trained models or ASR often rely on offline batch processing or incur prohibitive computational costs, rendering them unsuitable for streaming VC. Thus, online VC must balance efficiency and accuracy. Emformer’s chunk‐based incremental attention mechanism uses cached contextual information from previous chunks to maintain context continuity and capture long‐range dependencies, which preserves high content accuracy while substantially reducing latency, offering an efficient and practical solution for real‐time VC content extraction.
Our training objective is to distill the content representations extracted by HuBERT [27] from the source speech at 20 ms intervals—representations that have been widely adopted and validated in prior work [3]. As shown in Figure 1 (b), we base our design on Emformer by partitioning the input into chunks for both attention computation and subsequent network processing. During attention computation, we supply context keys, values, and memory‐bank embeddings extracted from the preceding chunk. The memory bank functions provide prior contextual information, with each encapsulating one chunk’s worth of information. At layer $n$ and chunk $i$, we compute:
$Q_{i}^{n}\;=\;\bigl{[}\,W_{q}\,C_{i}^{n},\;W_{q}\,R_{i}^{n},\;W_{q}\,s_{i}^{n}\bigr{]},$ (1)
$K_{i}^{n}\;=\;\bigl{[}\,W_{k}\,M_{i}^{n},\;K_{L,i}^{n},\;W_{k}\,C_{i}^{n},\;W_{k}\,R_{i}^{n}\bigr{]},$ (2)
$V_{i}^{n}\;=\;\bigl{[}\,W_{v}\,M_{i}^{n},\;V_{L,i}^{n},\;W_{v}\,C_{i}^{n},\;W_{v}\,R_{i}^{n}\bigr{]},$ (3)
$C_{i}^{\,n+1}\;=\;\mathrm{FFN}\Bigl{(}\mathrm{Attn}\bigl{(}W_{q}\,C_{i}^{n},\;K_{i}^{n},\;V_{i}^{n}\bigr{)}\;+\;C_{i}^{n}\Bigr{)},$ (4)
$m_{i}^{\,n+1}\;=\;\mathrm{Attn}\bigl{(}W_{q}\,s_{i}^{n},\;K_{i}^{n},\;V_{i}^{n}\bigr{)},$ (5)
$z_{c\text{-}i}\;=\;\arg\max_{1\leq s\leq S}\Bigl{[}\,\mathrm{Softmax}\bigl{(}U\,C_{i}^{(N)}\bigr{)}\Bigr{]}_{(s,t)}\;.$ (6)
Here, $C_{i}^{n}$ is the content‐chunk feature at layer $n,1<n<N$, $R_{i}^{n}$ is the right‐context block, $s_{i}^{n}$ is the chunk summary, and $M_{i-1}^{n}$ is the memory bank from the previous chunk. Attention is computed using $(Q_{i}^{n},K_{i}^{n},V_{i}^{n})$, the output is added residually to $C_{i}^{n}$, and fed into $\mathrm{FFN}$. The summary $m_{i}^{\,n+1}$ is obtained similarly. Finally, $z_{c\text{-}i}$ is produced by projecting $C_{i}^{(N)}$ with $U$, applying Softmax over $S$ classes per frame $t$, and selecting the highest‐probability label for each frame. All layers remain causal by caching only $M_{i-1}^{n}$, $K_{L,i}^{n}$, and $V_{L,i}^{n}$, enabling low‐latency, zero‐shot online VC content extraction.
In this attention mechanism, keys, values, and memory‐bank vectors are employed to capture contextual dependencies between data chunks. By flexibly adjusting the size of the right context, one can trade off performance against latency. Setting it to zero enforces a strictly causal behavior. The sizes of the left context and memory buffer determine the overall receptive field. Leveraging our Stream Content Extractor, we achieve efficient, high‐quality streaming content extraction.
III-C Adaptive Style Encoder
To comprehensively capture and transfer styles from reference speech, we introduce the Adaptive Style Encoder. As shown in Figure 2, reference mel-spectrogram $m_{rf}$ is initially refined through convolutional blocks. Subsequently, a downsampling layer pools the output into chunk-level representations. In our experiments, we set the chunk size to 80 ms to effectively incorporate emotion and prosody. Then, convolutional stacks capture correlations at the chunk level.
Next, we employ a linear projection to map the outputs into a low-dimensional latent space for efficient code index lookup [37]. The Clustering Vector Quantization (CVQ) layer [38] extracts style representations from these latent inputs, establishing an information bottleneck that effectively filters out non-style information [39]. Through dimensionality reduction via linear projection and the bottleneck mechanism of CVQ, we decouple speech styles from timbre and content.
Compared to traditional VQ [40], CVQ employs a dynamic initialization strategy during training. This ensures that less frequently or unused code vectors are updated more substantially than frequently used ones, effectively addressing the codebook collapse issue [38]. We train the CVQ layer using a combination of the VQ loss and contrastive loss, defined as:
$\mathcal{L}_{CVQ}=$ $\|sg[z_{e}(x)]-e\|^{2}_{2}+\beta\|z_{e}(x)-sg[e]\|^{2}_{2}+\mathcal{L}_{Contras},$ (7)
where $z_{e}(x)$ is the encoder output for input $x$, $e$ is the selected code from the codebook, $\text{sg}(\cdot)$ is the stop-gradient operator, and $\beta$ is commitment loss. The contrastive loss is:
$\mathcal{L}_{Contras}=-\log\frac{e^{\text{sim}(e_{k},\hat{z}_{i}^{+})/\tau}}{\sum*{i=1}^{N}e^{\text{sim}(e_{k},\hat{z}_{i}^{-})/\tau}}.$ (8)
For each code $e_{k}$, we directly select the closest $\hat{z}_{i}^{+}$ as the positive pair and sample other farther $\hat{z}_{i}^{-}$ as negative pairs.
Upon generating the style embedding, it is crucial to adaptively align this embedding with the sum $z_{ct-i}$ of $z_{c-i}$ and $z_{t}$ [41]. For this purpose, we introduce the Align Attention module, incorporating the Scaled Dot-Product Attention mechanism [42]. Before inputting the detailed style embedding into the attention module, we include positional encoding embeddings. Within this module, $z_{ct-i}$ serves as the query, while the style embedding functions as both key and value. Ultimately, this process yields aligned style $z_{s-i}$, effectively facilitating the modeling and transfer of reference styles.
Figure 2: The architecture of Adaptive Style Encoder.
III-D Causal Shuffle Vocoder
The chunkwise processing in online VC makes it incompatible with offline VC models and introduces boundary artifacts. Enforcing causality through autoregressive methods increases latency, while using zero-padding in non-causal vocoders leads to spectral artifacts. To address this, after generating the mel-spectrogram $m_{gn-i}$ using a causal mel decoder composed of causal convolutions, we build a fully causal streaming vocoder. Based on high-quality HiFi-GAN [43], we replace standard convolutions with causal convolutions and implement upsampling via pixel shuffle, resulting in a strictly causal vocoder with high audio quality, named Causal Shuffle Vocoder.
As illustrated in Figure 1 (c), $m_{gn-i}$ first passes through a single causal convolutional layer, producing an initial feature map of shape $(C,T_{0})$. The network then comprises $N$ upsampling stages. At the $n$-th stage, a causal convolution projects the channel dimension from $C$ up to $r_{n}C$, yielding an intermediate feature map of shape $(r_{n}C,T_{n-1})$. Next, pixel shuffling rearranges these $r_{n}C$ channels into a new tensor of shape $(C,r_{n}T_{n-1})$, $r_{n}T_{n-1}=T_{n}$, thereby performing temporal upsampling. Because pixel shuffling only reorders channels without introducing information from future frames, the network’s output at time step $t$ depends solely on input frames from time $1$ through $t$. By choosing the upsampling factors, the final temporal resolution becomes $T_{0}\prod_{n=1}^{N}r_{n}$, recovering the original audio sampling rate.
Immediately following each upsampling stage is a parallel residual block, which refines harmonic and transient details while preserving causality. Within each block, the input and its output are combined via a skip connection, allowing the network to learn only the incremental details rather than the entire spectral envelope. This design ensures fine-grained correction of the spectrum without ever accessing future frames. The outputs of all parallel residual branches at a given scale are averaged before being forwarded to the next upsampling layer. After the last upsampling layer and its associated residual block, a final causal convolution then maps the channel dimension to 1, producing the generated speech.
By combining causal convolution with pixel shuffling, the waveform produced at each frame depends exclusively on past mel-spectrogram samples. Causal convolutions pad solely on the left, avoiding spectral distortions from symmetric or right‐side padding. Moreover, pixel shuffling upscales by rearranging channels into a finer temporal grid without waiting for future frames, thus eliminating checkerboard artifacts. This approach matches HiFi‐GAN quality in an online, low‐latency setting because each step is a single forward pass, avoiding cumulative delay. Moreover, since pixel shuffling uniformly distributes low-resolution features along the time axis, the model can more effectively reconstruct audio details, yielding clear, natural speech even under strict causal constraints.
III-E Training and Inference Procedures
Figure 3: The chunk-wise online inference procedure of Conan. Here, we illustrate the causal setting without utilizing any right context.
Training Procedure. First, for Stream Content Extractor, we use the cross-entropy between the predicted content label and the label extracted by HuBERT. During training of the main model, we adopt the multi-length discriminator [44] to maintain perceptual quality. The total loss comprises $\mathcal{L}_{\mathrm{CVQ}}$ for the adaptive style encoder; $\mathcal{L}_{\mathrm{pitch}}$, the MSE between the ground truth and the predicted $F_{0}$; $\mathcal{L}_{\mathrm{mae}}$ and $\mathcal{L}_{\mathrm{ssim}}$, the MAE and SSIM losses between the predicted and ground truth mel‐spectrograms; and $\mathcal{L}_{\mathrm{GAN}}$, the LSGAN‐style adversarial loss [45], whose objective is to minimize the distribution distance between predicted and ground truth mel‐spectrograms. Finally, the loss for Causal Shuffle Vocoder follows HiFiGAN [43], including the GAN loss, mel loss, and feature‐map loss.
Inference Procedure. As illustrated in Figure 3, we first feed the entire reference speech into the model to provide timbre and stylistic information. During chunk‐wise online inference, we wait until the input reaches a predefined chunk size before passing it to the model, because our generation speed for each chunk is faster than the chunk’s duration, online generation becomes possible. To ensure temporal continuity, we employ a sliding context‐window strategy. Concretely, at each generation step, we not only input the source speech corresponding to the current chunk but also include the preceding context within the model’s receptive field. From the model’s output, we then extract only the segment required for this chunk, thus preserving causality. As a result, overlapping portions generated across successive chunks remain identical, yielding perfectly smooth boundaries between concatenated segments.
IV Experiments
IV-A Experimental Setup
Dataset. We use the LibriTTS dataset [46] for our experiments, which contains 586 hours of English speech from 2,456 speakers. All speech is downsampled to 16 kHz. To evaluate Conan’s performance in zero‐shot scenarios, we generalized the source model to the VCTK dataset, which comprises 44 hours of clean speech from 110 English speakers. For a fair comparison, we reference the StreamVC [3] data‐partitioning strategy. We use the test‐clean subset of LibriTTS as the source speech and then randomly select 6 unseen speakers from the VCTK dataset to form the out‐of‐domain test set.
Implementation Details. We set the sample rate to 16 kHz, the window size to 1024, the hop size to 320, and the number of mel bins to 80 to derive mel-spectrograms from raw waveforms. We set the chunk size to 80 ms and employ a 6-layer Emformer with 2 right context chunks as the full setting. For the fastest setting, the chunk size is 20 ms, and a 3-layer Emformer without right context chunks is employed. The default size of the codebook for CVQ is 128.
Training Details. We train our model using 1 NVIDIA A100 GPU. The Adam optimizer is used with $\beta_{1}=0.9$ and $\beta_{2}=0.98$. The Stream Content Extractor trains for 80k steps, the main model trains for 160k steps, and the Causal Shuffle Vocoder trains for 600k steps to reach convergence.
Evaluation Details. We objectively and subjectively evaluate our model based on three criteria: real‐time content accuracy, conversion quality, and speaker similarity. Content accuracy is measured by Word Error Rate (WER) and Character Error Rate (CER) using a HuBERT‐Large ASR model  [^1] https://github.com/resemble-ai/Resemblyzer. Speaker similarity (SIM) is assessed with Resemblyzer  [^2] https://huggingface.co/facebook/hubert-large-ls960-ft. For subjective metrics, we conduct mean opinion score (MOS) and comparative mean opinion score (CMOS) evaluations. Both metrics are rated on a 1–5 scale and reported with 95% confidence intervals. Specifically, we use MOS‐Q to judge conversion quality (for clarity and naturalness) and MOS‐S to assess speaker similarity (for timbre and styles) [47]. For CMOS-Q and CMOS-S, listeners are tasked with comparing pairs of speech produced by different systems and expressing their preferences. All participants are fairly compensated and agree that the results will be used for scientific research.
IV-B Main Results
TABLE I: Evaluation results of Conan in comparison with baseline approaches. We conduct both subjective and objective evaluations.
Method Method Content Accuracy Speaker Similarity Conversion Quality
WER $\downarrow$ CER $\downarrow$ SIM $\uparrow$ MOS-S $\uparrow$ MOS-Q $\uparrow$
Data Source 5.41% 2.43% - - 3.96 $\pm$ 0.08
Source (Vocoder) 5.56% 2.67% - - 3.94 $\pm$ 0.15
Reference (Vocoder) - - 95.81% 4.43 $\pm$ 0.14 -
Baseline VQMIVC [4] 56.73% 36.18% 62.54% 3.76 $\pm$ 0.05 3.81 $\pm$ 0.07
BNE-PPG-VC [48] 12.29% 4.87% 79.06% 3.90 $\pm$ 0.13 3.86 $\pm$ 0.14
Diff-VCTK [49] 10.58% 5.43% 81.47% 3.95 $\pm$ 0.07 4.05 $\pm$ 0.04
QuickVC [50] 7.14% 3.10% 77.39% 3.84 $\pm$ 0.04 3.99 $\pm$ 0.09
StreamVC [3] 6.22% 2.17% 77.81% 3.86 $\pm$ 0.05 4.01 $\pm$ 0.13
Ours Conan (Full) 6.02% 2.83% 85.71% 4.02 $\pm$ 0.06 4.06 $\pm$ 0.04
Conan (Fastest) 10.72% 4.74% 78.26% 3.85 $\pm$ 0.13 3.99 $\pm$ 0.11
We use several popular voice‐conversion models, including VQMIVC [4], BNE‐PPG‐VC [48], Diff‐VCTK [49], QuickVC [50], and StreamVC [3]. Since StreamVC is not open source, to ensure a fair comparison, we adopted the same dataset split and objective evaluation protocols they used. We rely on the objective metrics reported in their paper and perform subjective assessments based on their demo pages. Specifically, we collect utterances from speakers in the test-clean subset of LibriTTS as the source speech, and randomly select 6 speakers from the VCTK dataset as the reference speech sample.
As shown in Table I, Conan in the full setting delivers the highest speaker similarity (SIM, MOS-S), indicating that the Adaptive Style Encoder effectively models the speaker’s style and therefore achieves superior zero-shot similarity. Conan also attains the best conversion quality (MOS-Q), demonstrating that the streaming framework enhances both synthesis quality and naturalness. For content accuracy, Conan records the lowest WER, while its CER is slightly higher than that of StreamVC. This is mainly because StreamVC does not attempt style transfer and instead reuses the source pitch and energy distributions, giving it a CER advantage. Finally, the metrics for Source (Vocoder) and Reference (Vocoder) are both excellent, further highlighting the outstanding online reconstruction capability of the Causal Shuffle Vocoder.
TABLE II: Latency and RTF for Conan’s full and fastest settings tested on a single A100 80 GB GPU. Latency includes the sum of each module’s delay, plus the chunk size, and the right context.
Method Full Setting Fastest Setting
RTF Latency (ms) RTF Latency (ms)
Content 0.07 5.60 0.14 2.76
Main 0.10 7.88 0.39 7.82
Vocoder 0.08 6.21 0.31 6.29
Overall 0.25 139.71 0.74 36.87
As shown in Table II, in the faster setting, Conan uses a 20 ms chunk size to perform strictly causal, chunk‐wise online voice conversion with only 37 ms of latency, yet still achieves high performance metrics. Under Conan’s full setting with 80 ms chunk size and 2 right context chunks, we achieve substantially higher quality, nearly matching offline VC performance, while still meeting real‐time latency constraints.
IV-C Ablation Study
TABLE III: Evaluation results for the ablation study. SCE denotes Stream Content Encoder, ASE is Adaptive Style Encoder, and CSV represents Causal Shuffle Vocoder.
[…]
Conan 6.02% 0.00 0.00
w/o SCE 26.92% -0.33 -0.18
w/o Right 8.57% -0.19 -0.07
w/o ASE 6.53% -0.17 -0.19
w/o CSV 7.06% -0.21 -0.05
As shown in Table III, we conduct ablation studies to demonstrate the effectiveness of various design choices within Conan. When the original HuBERT is used instead of the Stream Content Encoder, we observe a substantial deterioration in both synthesis quality and WER, underscoring the incompatibility of HuBERT with streaming scenarios. Furthermore, when right‐context chunks are omitted, both MOS‐Q and WER worsen, indicating that the size of the lookahead can be adjusted as a trade‐off between efficiency and quality. Furthermore, removing the Adaptive Style Encoder results in a marked decrease in MOS-S, confirming that our approach effectively captures speaker styles. Finally, substituting the Causal Shuffle Vocoder with a zero-padded causal HiFi-GAN leads to a decline in MOS-Q, as the zero padding introduces checkerboard artifacts, causing a decline in synthesis quality.
V Conclusion
In this paper, we introduce Conan, a method that converts speech by preserving the original content while adopting the timbre and styles of a reference utterance. At its core, Conan employs the Streaming Content Extractor based on Emformer to encode source audio with minimal latency, ensuring semantic information is captured in real time. To adapt to styles of reference speech, Adaptive Style Encoder analyzes the reference speech and derives detailed stylistic representations that guide the conversion process. Finally, Causal Shuffle Vocoder, built upon a fully causal HiFiGAN architecture enhanced with pixel-shuffle operations, reconstructs high-fidelity audio with low generation delay. Experimental results indicate that Conan surpasses existing baseline approaches, achieving higher scores in both subjective and objective metrics.